# Distribution of local signs of modular forms and murmurations of Fourier coefficients

### v2

@@ Abstract.
Recently, we showed that global root numbers of modular forms are biased toward $+1$. Together with Pharis, we also showed an initial bias of Fourier coefficients towards the sign of the root number. First, we prove analogous results with respect to local root numbers.
@@ 1.1. Distributions of local root numbers
In [me:refdim, me:rootno], we observed a bias of newforms towards global root number $+1$, even though asymptotically these account for 50% of newforms. Namely, for fixed pair $(k,N)$, outside of a prescribed set of exceptions, there are strictly more newforms in $S^{\mathrm{new}}_{k}(N)$ with root number $+1$ than $-1$. Moreover, the excess number of forms with root number $+1$ is essentially independent of $k$, and is typically an elementary factor times the class number of $\mathbb{Q}(\sqrt{-N})$.
In fact [me:refdim] was primarily concerned with the distribution of local root numbers in $S_{k}^{\mathrm{new}}(N)$. Suppose $N$ is squarefree, and $q_{1},\dots,q_{m}$ are primes dividing $N$. We obtained a criterion for when the local root numbers (i.e., Atkin–Lehner eigenvalues) at $q_{1},\dots,q_{m}$ are perfectly equidistributed in $S_{k}^{\mathrm{new}}(N)$, i.e., the number of newforms with prescribed local signs at $q_{1},\dots,q_{m}$ does not depend on the choice of signs. We also showed that there is a bias towards/away all local signs being $-1$, with the direction of the bias depending on the parities of $\frac{k}{2}$ and the number of prime divisors of $N$.
The motivation for studying the distribution of local root numbers in [me:refdim] was for applications to congruences mod 2. Suppose further that $N$ is a squarefree product of an odd number of primes. In [me:cong2], we showed that, apart from levels of the form $N=2p_{1}p_{2}$ when $k=2$, if the local Atkin–Lehner signs are perfectly equidistributed for $q_{1},\dots,q_{m}$, then for any newform $f\in S_{k}(N)$ and any prescribed choice of local signs at $q_{1},\dots,q_{m}$, there is a newform $g\in S_{k}(N)$ with those prescribed signs which is congruent to $f$ mod $2$. Further when $k=2$, where there is a bias towards all local signs $-1$, one can do the same (without perfect equidistribution) when one prescribes the local signs of $g$ to all be $-1$.
@@
As a consequence, outside of certain exceptional cases, we characterize when $\Delta_{k}(q^{r},M)=0$, and specify the sign of $\Delta_{k}(q^{r},M)$ when it is nonzero. Let $\omega(M)$ be the number of primes dividing $M$, $\omega_{1}(M)$ the number of primes sharply dividing $M$, and $\omega_{2}(n;M)$ the number of $p^{2}\parallel M$ such that (${n\overwithdelims( )p}=1$.
@@
Then $\Delta_{k}(q^{r},M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) (${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $e\geq 5$; (iv) $e=4$ and $q\equiv 1\bmod 4$; or (v) $e=1,2,3$ and $q\equiv 7\bmod 8$.
Moreover, when $\Delta_{k}(q^{r},M)\neq 0$, its sign is $(-1)^{k/2+\omega_{1}(M^{\prime})+\omega_{2}(-q,M^{\prime})}b_{q,e}$, where $b_{q,e}$ is the sign of the quantity $\alpha_{1}(-q;e)$ in Table 2. I.e., we may take $b_{q,e}=+1$ if $e=0$; $-1$ if $e=1,2$; (${-1\overwithdelims( )q}$ if $e=3$; and ($-{2\overwithdelims( )q}$ if $e=4$. Hence this sign is simply $(-1)^{k/2+\omega(M)}$ when $M$ is squarefree.
@@
Due to the $\delta_{r=2}(k-1)D_{2}$ term in (1.1), the behavior is different when $r=2$. In this case we describe the asymptotic behavior, which in the following setting asserts a bias towards local root number $-1$.
@@
We also establish the asymptotic behavior in $q$ under local conditions on $M$ (see 4.4). The fact that local root number distributions behave differently in $r=2$ parallels the fact that the bias of global root numbers is different for levels which are perfect squares (see [me:rootno]). We do not have a compelling intuitive explanation for why this is (for local or global root numbers), but we do note that, when $q$ is odd, $r=2$ is precisely the case where the class of possible local representations $\pi_{q}$ at $q$ associated to newforms $f\in S_{k}^{\mathrm{new}}(q^{r}N)$ includes ramified principal series and ramified twists of Steinberg representations, all of which have local root number (${-1\overwithdelims( )q}$. However, such forms cannot account for the bias towards local root number $-1$ when $q\equiv 1\bmod 4$.
@@ 1.2. Correlation of initial Fourier coefficients with local signs
In [me:pharis], we showed that for squarefree levels $N$, the trace of a Hecke operator $T_{\ell}$ on the subspace of forms in $S_{k}^{\mathrm{new}}(N)$ with root number $+1$ (resp., $-1$) is positive (resp., negative) for $\ell$ small relative to $N$ [^1] There is a sign error for this result in [me:pharis] when $k\equiv 0\bmod 4$. We correct this in Appendix A. In other words, for small $\ell$, the sign of the Fourier coefficients $a_{\ell}(f)$ are biased towards the sign of the root number of $f$. Combined with [me:refdim], this means that for small $\ell$ the Fourier coefficients $a_{\ell}(f)$ have a positive (resp., negative) bias for forms with the more common (resp., less common) global root number.
@@
    Either $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}$ is 0 or it has the same sign as $\Delta_{k}(q,M)$. It is 0 if and only if (i) (${-q\ell\overwithdelims( )p}=1$ for some odd $p\mid M$ or (ii) $M$ is even and $q\ell\equiv 7\bmod 8$.
@@
    If $q$ is sufficiently large with respect to $k,M,\ell$, and if $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}\neq 0$, then the trace of $T_{\ell}$ on $S_{k}^{\mathrm{new}}(qM)^{\pm_{q}}$ has the same sign as ±$\pm\Delta_{k}(q,M)$.
@@ 1.3. Murmurations
Fix $k\geq 2$ even and let $\mathcal{F}=\mathcal{F}_{k}$ be a suitably large family of weight $k$ newforms, say all weight $k$ newforms (with trivial nebentypus), or all of those with squarefree level. Let $\mathcal{F}(X,Y)$ (resp. ±$\mathcal{F}^{\pm}(X,Y)$) be the set of newforms $f\in\mathcal{F}$ with level $X\leq N\leq Y$ (resp. and have have root number ±$\pm 1$). Denote by $\mathcal{F}(X,Y)^{(\ell)}$ (resp. ±$\mathcal{F}^{\pm}(X,Y)^{(\ell)}$ be the subset of such forms with level $N$ coprime to $\ell$. The murmuration phenomena is the numerical observation that, for a fixed $\beta>1$, the averages
@@
tend to continuous murmuration functions $M^{\pm}_{\mathcal{F}}(x;\beta)$ as $\ell,X\to\infty$ such that $\frac{\ell}{X}\to x$. The restriction to forms in ±$\mathcal{F}^{\pm}(X,Y)^{(\ell)}$, as opposed to ±$\mathcal{F}^{\pm}(X,Y)$, is just a computational convenience and will not affect asymptotics (see Section 6.4). The normalization factor $\ell^{1-\frac{k}{2}}$ weights the Fourier coefficients so that each term is $O(\sqrt{\ell})$. In all numerical calculations, we follow Sutherland and take $\beta=2$.
@@
See Sections 1.3 and 1.3 for plots of $A_{\mathcal{F}}^{\pm}(\ell,X;2)$ for $X=1000$ and $X=2000$ for the weight 2 squarefree level family. Blue dots represent root number $+1$ and red dots $-1$. We labeled the horizontal axes by $\ell$, but really one should think of the horizontal axes as being labeled by $\frac{\ell}{X}$, so both graphs have horizontal range $40<\frac{\ell}{X}<$ 4. That the graphs have a limit in this scale is called scale invariance in $\frac{\ell}{X}$. The limiting murmuration functions $M^{\pm}_{\mathcal{F}}$ oscillate infinitely [zubrilina].
@@
Zubrilina [zubrilina] proved such murmurations for $A_{\mathcal{F}}$ when $\mathcal{F}$ is the family of weight $k$ newforms of squarefree level. In fact, Zubrilina proves a localized version of murmurations—essentially this means one can work with intervals of the form $[X,\beta X]$ where $\beta\to 0$ as $X\to\infty$. This leads to a continuous murmuration density function, and one obtains the murmuration functions $M_{\mathcal{F}}(x;\beta)$ by integrating the murmuration density function (and this implies continuity in $\beta$).
@@
where $\mathcal{Q}$ is a sequence of divisors $Q\mid N$ for each level $N$ appearing the family $\mathcal{F}$. Here $\mathcal{F}(N)=\mathcal{F}(N,N)$, and $w_{Q}(f)$ is the $W_{Q}$-eigenvalue of $f$. The notation $\sum^{\prime}$ means we restrict to summing $N$ coprime to $\ell$. The normalization factor $\sqrt{\frac{N}{Q}}$ is included to keep the averages at about the same size so they do not tend to 00 as $X\to\infty$ (see Section 6.2).
Note that if each $Q=N$ then (1.3) becomes (1.2). At the other extreme if each $Q=1$ then we are considering sums without any root numbers (as $w_{1}(f)=1$). See Sections 1.3 and 1.3 for plots of the averages in (1.3) in this case when $k=2$. (Graphs for $k=4$ are roughly similar.) Without the normalization factor of $\sqrt{\frac{N}{Q}}=\sqrt{N}$ in this case, the analogues of Sections 1.3 and 1.3 are graphs which individually have similar shapes, but whose vertical scale shrinks with $X$. That the unweighted averages for $Q=1$ tend to 00 reflects the root number symmetry $M^{+}_{\mathcal{F}}(x;\beta)=-M^{-}_{\mathcal{F}}(x;\beta)$.
@@
    for a murmuration function $M_{\mathcal{F}}^{\mathcal{Q}}$ which is continuous on $[0,\infty)\times(1,\infty)$.
@@
    for a murmuration function $\tilde{M}_{\mathcal{F}}^{\mathcal{Q},\delta}$ which is continuous on $[0,\infty)\times(1,\infty)$.
Recall that Type I murmurations with $N=Q$ squarefree and $k=2$ are illustrated in Sections 1.3 and 1.3. See Section 1.3 for a Type I plot with $N=5Q$ ($Q$ squarefree) and $k=4$, averaging over the range $3000015000<N<$ 30000. Similarly, non-smoothed and smoothed Type II plots with $Q=1$ are given in Sections 1.3, 1.3, 1.3 and 1.3. Illustrations of Type III situations with $N=pq$ are presented in a different format in Sections 1.3, 1.3 and 1.3 by looking at Atkin–Lehner eigenspaces, as will be described below.
@@
If one works with (squarefree or general) levels $N$ that have a fixed number of prime divisors as in Type III, then one can alternatively look how the collection of Atkin–Lehner signs is correlated with Fourier coefficients. See Sections 1.3 and 1.3 for a graph of averages of $a_{\ell}$’s over newforms in $S_{2}(pq)$ and $S_{4}(pq)$ with fixed Atkin–Lehner signs at $p,q$, where $p=2$ and $60003000<q<$ 6000. The blue and green dots correspond to signs ++ and -- and red and orange dots to signs +- and -+, respectively. (The first sign denotes the sign at $p$, and the second the sign at $q$.) See Section 1.3 for the analogous graph for $S_{4}(pq)$ where $p,q$ both vary such that $p<q$ and $120006000<pq<$ 12000.
@@
for a murmuration function $M_{\mathcal{F}}^{\varepsilon}$ which is continuous on $[0,\infty)\times(1,\infty)$.
@@
Figure 11. $w_{Q}$ (blue) versus $w_{N}$ (red) murmurations for elliptic curves of squarefree conductor $N=2Q$, with $4000020000<Q<$ 40000
@@ 3. Traces on newspaces
(Throughout the analysis in this section, $q,\ell,k$ will be fixed, so we suppress them from our notation for brevity.) For $r<0$, we interpret these quantities to be 00. When $\ell=1$, $t^{\mathrm{new}}(r;M)=\Delta_{k}(q^{r},M)$.
@@
$t(r;M)-t(r-2;M)=\begin{cases}A_{1,0}(r;M)+A_{2}(r;M)+A_{3}&\text{if }r=0,1,\\ A_{1,0}(r;M)-A_{1,0}(r-2;M)-A_{1,1}(r-2;M)\\ \quad{}+\delta_{r\geq 4}A_{1,1}(r-4;M)+A_{2}(r;M)-A_{2}(r-2;M)&\text{if }r\geq 2.\\ \end{cases}$
@@
(3.3) $t^{\mathrm{new}}(r;M)=\begin{cases}\tilde{A}_{1,0}(r;M)+\tilde{A}_{2}(r;M)+\tilde{A}_{3}&\text{if }r=0,1,\\ \tilde{A}_{1,0}(r;M)-\tilde{A}_{1,0}(r-2;M)-\tilde{A}_{1,1}(r-2;M)\\ \quad{}+\delta_{r\geq 4}\tilde{A}_{1,1}(r-4;M)+\tilde{A}_{2}(r;M)-\tilde{A}_{2}(r-2;M)&\text{if }r\geq 2.\\ \end{cases}$

### v3

@@ Abstract.
Recently, we showed that global root numbers of modular forms are biased toward +1+1. Together with Pharis, we also showed an initial bias of Fourier coefficients towards the sign of the root number. First, we prove analogous results with respect to local root numbers.
@@
Corrected version (October 30, 2025) [^1] After this paper appeared in Mathematika (2025), I discovered minor mathematical misprints in Section 4.2. This version corrects those errors. (See footnote in Section 4.2 for changes.)
@@ 1.1. Distributions of local root numbers
In [me:refdim, me:rootno], we observed a bias of newforms towards global root number +1+1, even though asymptotically these account for 50% of newforms. Namely, for fixed pair $(k,N)$, outside of a prescribed set of exceptions, there are strictly more newforms in $S^{\mathrm{new}}_{k}(N)$ with root number +1+1 than 1-1. Moreover, the excess number of forms with root number +1+1 is essentially independent of $k$, and is typically an elementary factor times the class number of $\mathbb{Q}(\sqrt{-N})$.
In fact [me:refdim] was primarily concerned with the distribution of local root numbers in $S_{k}^{\mathrm{new}}(N)$. Suppose $N$ is squarefree, and $q_{1},\dots,q_{m}$ are primes dividing $N$. We obtained a criterion for when the local root numbers (i.e., Atkin–Lehner eigenvalues) at $q_{1},\dots,q_{m}$ are perfectly equidistributed in $S_{k}^{\mathrm{new}}(N)$, i.e., the number of newforms with prescribed local signs at $q_{1},\dots,q_{m}$ does not depend on the choice of signs. We also showed that there is a bias towards/away all local signs being 1-1, with the direction of the bias depending on the parities of $\frac{k}{2}$ and the number of prime divisors of $N$.
The motivation for studying the distribution of local root numbers in [me:refdim] was for applications to congruences mod 2. Suppose further that $N$ is a squarefree product of an odd number of primes. In [me:cong2], we showed that, apart from levels of the form $N=2p_{1}p_{2}$ when $k=2$, if the local Atkin–Lehner signs are perfectly equidistributed for $q_{1},\dots,q_{m}$, then for any newform $f\in S_{k}(N)$ and any prescribed choice of local signs at $q_{1},\dots,q_{m}$, there is a newform $g\in S_{k}(N)$ with those prescribed signs which is congruent to $f$ mod $2$. Further when $k=2$, where there is a bias towards all local signs 1-1, one can do the same (without perfect equidistribution) when one prescribes the local signs of $g$ to all be 1-1.
@@
As a consequence, outside of certain exceptional cases, we characterize when $\Delta_{k}(q^{r},M)=0$, and specify the sign of $\Delta_{k}(q^{r},M)$ when it is nonzero. Let $\omega(M)$ be the number of primes dividing $M$, $\omega_{1}(M)$ the number of primes sharply dividing $M$, and $\omega_{2}(n;M)$ the number of $p^{2}\parallel M$ such that ${n\overwithdelims( )p}=1$.
@@
Then $\Delta_{k}(q^{r},M)=0$ if and only if (i) $M^{\prime}$ is not cubefree; (ii) ${-q\overwithdelims( )p}=1$ for some $p\parallel M^{\prime}$; (iii) $e\geq 5$; (iv) $e=4$ and $q\equiv 1\bmod 4$; or (v) $e=1,2,3$ and $q\equiv 7\bmod 8$.
Moreover, when $\Delta_{k}(q^{r},M)\neq 0$, its sign is $(-1)^{k/2+\omega_{1}(M^{\prime})+\omega_{2}(-q,M^{\prime})}b_{q,e}$, where $b_{q,e}$ is the sign of the quantity $\alpha_{1}(-q;e)$ in Table 2. I.e., we may take $b_{q,e}=+1$ if $e=0$; 1-1 if $e=1,2$; $1\overwithdelims( )q$} if $e=3$; and $2\overwithdelims( )q$} if $e=4$. Hence this sign is simply $(-1)^{k/2+\omega(M)}$ when $M$ is squarefree.
@@
Due to the $\delta_{r=2}(k-1)D_{2}$ term in (1.1), the behavior is different when $r=2$. In this case we describe the asymptotic behavior, which in the following setting asserts a bias towards local root number 1-1.
@@
We also establish the asymptotic behavior in $q$ under local conditions on $M$ (see 4.4). The fact that local root number distributions behave differently in $r=2$ parallels the fact that the bias of global root numbers is different for levels which are perfect squares (see [me:rootno]). We do not have a compelling intuitive explanation for why this is (for local or global root numbers), but we do note that, when $q$ is odd, $r=2$ is precisely the case where the class of possible local representations $\pi_{q}$ at $q$ associated to newforms $f\in S_{k}^{\mathrm{new}}(q^{r}N)$ includes ramified principal series and ramified twists of Steinberg representations, all of which have local root number $1\overwithdelims( )q$}. However, such forms cannot account for the bias towards local root number 1-1 when $q\equiv 1\bmod 4$.
@@ 1.2. Correlation of initial Fourier coefficients with local signs
In [me:pharis], we showed that for squarefree levels $N$, the trace of a Hecke operator $T_{\ell}$ on the subspace of forms in $S_{k}^{\mathrm{new}}(N)$ with root number +1+1 (resp., 1-1) is positive (resp., negative) for $\ell$ small relative to NN. [^2] There is a sign error for this result in [me:pharis] when $k\equiv 0\bmod 4$. We correct this in Appendix A. In other words, for small $\ell$, the sign of the Fourier coefficients $a_{\ell}(f)$ are biased towards the sign of the root number of $f$. Combined with [me:refdim], this means that for small $\ell$ the Fourier coefficients $a_{\ell}(f)$ have a positive (resp., negative) bias for forms with the more common (resp., less common) global root number.
@@
    Either $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}$ is 0 or it has the same sign as $\Delta_{k}(q,M)$. It is 0 if and only if (i) ${-q\ell\overwithdelims( )p}=1$ for some odd $p\mid M$ or (ii) $M$ is even and $q\ell\equiv 7\bmod 8$.
@@
    If $q$ is sufficiently large with respect to $k,M,\ell$, and if $\operatorname{tr}_{S_{k}^{\mathrm{new}}(qM)}T_{\ell}W_{q}\neq 0$, then the trace of $T_{\ell}$ on $S_{k}^{\mathrm{new}}(qM)^{\pm_{q}}$ has the same sign as $\pm\Delta_{k}(q,M)$.
@@ 1.3. Murmurations
Fix $k\geq 2$ even and let $\mathcal{F}=\mathcal{F}_{k}$ be a suitably large family of weight $k$ newforms, say all weight $k$ newforms (with trivial nebentypus), or all of those with squarefree level. Let $\mathcal{F}(X,Y)$ (resp. $\mathcal{F}^{\pm}(X,Y)$) be the set of newforms $f\in\mathcal{F}$ with level $X\leq N\leq Y$ (resp. and have have root number $\pm 1$). Denote by $\mathcal{F}(X,Y)^{(\ell)}$ (resp. $\mathcal{F}^{\pm}(X,Y)^{(\ell)}$ be the subset of such forms with level $N$ coprime to $\ell$. The murmuration phenomena is the numerical observation that, for a fixed $\beta>1$, the averages
@@
tend to continuous murmuration functions $M^{\pm}_{\mathcal{F}}(x;\beta)$ as $\ell,X\to\infty$ such that $\frac{\ell}{X}\to x$. The restriction to forms in $\mathcal{F}^{\pm}(X,Y)^{(\ell)}$, as opposed to $\mathcal{F}^{\pm}(X,Y)$, is just a computational convenience and will not affect asymptotics (see Section 6.4). The normalization factor $\ell^{1-\frac{k}{2}}$ weights the Fourier coefficients so that each term is $O(\sqrt{\ell})$. In all numerical calculations, we follow Sutherland and take $\beta=2$.
@@
See Sections 1.3 and 1.3 for plots of $A_{\mathcal{F}}^{\pm}(\ell,X;2)$ for $X=1000$ and $X=2000$ for the weight 2 squarefree level family. Blue dots represent root number +1+1 and red dots 1-1. We labeled the horizontal axes by $\ell$, but really one should think of the horizontal axes as being labeled by $\frac{\ell}{X}$, so both graphs have horizontal range $0<\frac{\ell}{X}<4$. That the graphs have a limit in this scale is called scale invariance in $\frac{\ell}{X}$. The limiting murmuration functions $M^{\pm}_{\mathcal{F}}$ oscillate infinitely [zubrilina].
@@
Zubrilina [zubrilina] proved such murmurations for $A_{\mathcal{F}}$ when $\mathcal{F}$ is the family of weight $k$ newforms of squarefree level. In fact, Zubrilina proves a localized version of murmurations—essentially this means one can work with intervals of the form $X,\beta X]$ where $\beta\to 0$ as $X\to\infty$. This leads to a continuous murmuration density function, and one obtains the murmuration functions $M_{\mathcal{F}}(x;\beta)$ by integrating the murmuration density function (and this implies continuity in $\beta$).
@@
where $\mathcal{Q}$ is a sequence of divisors $Q\mid N$ for each level $N$ appearing the family $\mathcal{F}$. Here $\mathcal{F}(N)=\mathcal{F}(N,N)$, and $w_{Q}(f)$ is the $W_{Q}$-eigenvalue of $f$. The notation $\sum^{\prime}$ means we restrict to summing $N$ coprime to $\ell$. The normalization factor $\sqrt{\frac{N}{Q}}$ is included to keep the averages at about the same size so they do not tend to 0 as $X\to\infty$ (see Section 6.2).
Note that if each $Q=N$ then (1.3) becomes (1.2). At the other extreme if each $Q=1$ then we are considering sums without any root numbers (as $w_{1}(f)=1$). See Sections 1.3 and 1.3 for plots of the averages in (1.3) in this case when $k=2$. (Graphs for $k=4$ are roughly similar.) Without the normalization factor of $\sqrt{\frac{N}{Q}}=\sqrt{N}$ in this case, the analogues of Sections 1.3 and 1.3 are graphs which individually have similar shapes, but whose vertical scale shrinks with $X$. That the unweighted averages for $Q=1$ tend to 0 reflects the root number symmetry $M^{+}_{\mathcal{F}}(x;\beta)=-M^{-}_{\mathcal{F}}(x;\beta)$.
@@
    for a murmuration function $M_{\mathcal{F}}^{\mathcal{Q}}$ which is continuous on $0,\infty)\times(1,\infty)$.
@@
    for a murmuration function $\tilde{M}_{\mathcal{F}}^{\mathcal{Q},\delta}$ which is continuous on $0,\infty)\times(1,\infty)$.
Recall that Type I murmurations with $N=Q$ squarefree and $k=2$ are illustrated in Sections 1.3 and 1.3. See Section 1.3 for a Type I plot with $N=5Q$ ($Q$ squarefree) and $k=4$, averaging over the range $15000<N<30000$. Similarly, non-smoothed and smoothed Type II plots with $Q=1$ are given in Sections 1.3, 1.3, 1.3 and 1.3. Illustrations of Type III situations with $N=pq$ are presented in a different format in Sections 1.3, 1.3 and 1.3 by looking at Atkin–Lehner eigenspaces, as will be described below.
@@
If one works with (squarefree or general) levels $N$ that have a fixed number of prime divisors as in Type III, then one can alternatively look how the collection of Atkin–Lehner signs is correlated with Fourier coefficients. See Sections 1.3 and 1.3 for a graph of averages of $a_{\ell}$’s over newforms in $S_{2}(pq)$ and $S_{4}(pq)$ with fixed Atkin–Lehner signs at $p,q$, where $p=2$ and $3000<q<6000$. The blue and green dots correspond to signs ++ and -- and red and orange dots to signs +- and -+, respectively. (The first sign denotes the sign at $p$, and the second the sign at $q$.) See Section 1.3 for the analogous graph for $S_{4}(pq)$ where $p,q$ both vary such that $p<q$ and $6000<pq<12000$.
@@
for a murmuration function $M_{\mathcal{F}}^{\varepsilon}$ which is continuous on $0,\infty)\times(1,\infty)$.
@@
Figure 11. $w_{Q}$ (blue) versus $w_{N}$ (red) murmurations for elliptic curves of squarefree conductor $N=2Q$, with $20000<Q<40000$
@@ 3. Traces on newspaces
(Throughout the analysis in this section, $q,\ell,k$ will be fixed, so we suppress them from our notation for brevity.) For $r<0$, we interpret these quantities to be 0. When $\ell=1$, $t^{\mathrm{new}}(r;M)=\Delta_{k}(q^{r},M)$.
@@
[…]
@@
[…]